A right triangle has a hypotenuse c=12.6 and a leg b=8.5. What does leg a approximately equal?
Answer by rfer(10417)

c^2=a^2+b^2
12.6^2-8.5^2=a^2
158.76-72.25=86.51
sq rt of 86.51=9.3